# Stable Neural Stochastic Differential Equations in Analyzing Irregular Time Series Data

### v1

@@ 2 Related Work & Preliminaries
Let $(\Omega,{\mathcal{F}},P)$ be a probability space. Fix integer $d\geq 1$. For an ${\mathbb{R}}^{d}$-valued random variable $X$, its law is denoted by ${\mathcal{L}}(X)$. ||$\cdot$|| ⋅ | is the Euclidean norm where the dimension of the space may vary depending on the context. We denote by $L^{p}(\Omega)$ the set of $X$ with $(\mathbb{E}[|X|^{p}])^{1/p}<\infty$. Let ${\mathcal{P}}({\mathbb{R}}^{d})$ denote the set of probability measures on ${\mathcal{B}}({\mathbb{R}}^{d})$. For two probability measures $\mu,\nu\in{\mathcal{P}}({\mathbb{R}}^{d})$, let ${\mathcal{C}}(\mu,\nu)$ denote the set of probability measures $\Pi$ on ${\mathcal{B}}({\mathbb{R}}^{2d})$ such that its respective marginals are $\mu$ and $\nu$. Then, the Wasserstein distance of order $p\geq 1$ is defined as
@@ 3.1 The proposed Neural SDEs
where the initial condition $h(\cdot;\theta_{h})$ is an affine function with parameter $\theta_{h}$, the drift function $\gamma(\cdot,\cdot;\theta_{\gamma})$ is a neural network with parameter $\theta_{\gamma}$ and the diffusion function $\sigma(\cdot;\theta_{\sigma})$ is a neural network with parameter $\theta_{\sigma}$. Due to its exponential form, Neural GSDE has distinct characteristics associated with deep ReLU networks such as a unique positive solution and an absorbing state of 00, which are not observed in Neural LSDE and Neural LNSDE (see Proposition 3.4 (ii)).
@@ 3.2 Properties of the proposed Neural SDEs
    Assume that the activation function in $\gamma$ is either a tanh or sigmoid function. Then, the SDE in Equation (6) has a unique strong solution ${\bm{z}}(t)$, which is nonnegative almost surely, i.e., $P({\bm{z}}(t)\geq 0\mbox{ for all $t\geq 0$})=1$ a.s. Furthermore, state 00 is an absorbing state.
The proof for Proposition 3.4 can be found in Appendix A. Intuitively, Proposition 3.4 (ii) implies that the latent representation ${\bm{z}}(t)$ of the Neural GSDE has always nonnegative values. Furthermore, once the one of values of the latent representation reaches 00, it remains there forever. This property is consistent with that of ReLU networks which have positive values for activated neurons and zeros for deactivated neurons. In particular, the fact that deactivated neurons never turn to activated neurons corresponds to the property of state 00 being an absorbing state.
@@ 3.3 Robustness under distribution shift
Let ${\bm{x}}$ denote the input data, and its law be represented by ${\mathcal{L}}({\bm{x}})$, i.e., ${\bm{x}}$ is a ${\mathbb{R}}^{d_{x}}$-valued random variable. Consider $\widetilde{\bm{x}}$ to be a perturbed version of the input data, with the law ${\mathcal{L}}(\widetilde{\bm{x}})$, such that
@@
${\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $\leq\sqrt{3}L_{F}L_{h}c_{1}e^{-c_{2}T}\sqrt{\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)}\rho,$
${\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $\leq\sqrt{2\sqrt{3}L_{h}c_{1}}L_{F}e^{-c_{2}T/2}\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)^{1/4}\sqrt{\rho},$
@@ Results.
Methods Mean Squared Error (×$\times 10^{-3}$)
@@
Table 3: Accuracy and memory usage (in MB) on Speech Commands
@@
Table 3: Accuracy and memory usage (in MB) on Speech Commands
@@ Acknowledgement
where $\gamma_{i}$ is the $i$-th element of $\gamma$ and $(\sigma(t;\theta_{\sigma}))_{ij}$ represents the $(i,j)$-th element of $\sigma(t;\theta_{\sigma})$, and $[{\bm{y}}_{i}(t),{\bm{y}}_{i}(t)]$ is the quadratic variation of ${\bm{y}}_{i}(t)$ up to $t$ for $i=1,2,\ldots,d_{z}$. In the matrix form, ${\bm{z}}(t)$ satisfies
@@
Therefore, ${\bm{z}}(t)$ is the solution of Equation (6). Let $T_{0}^{i}=\inf\{t>0|z_{i}(t)=0\}$ be the first hitting time to state 00 with $z_{i}(0)>0$ for $i=1,2,\ldots,d_{z}$ where $z_{i}(t)$ is the $i$-th component of ${\bm{z}}(t)$. Since ${\mathrm{d}}z_{i}(t)=0$ for $t\geq T_{0}^{i}$, then $z_{i}(t)=0$ for all $t\geq T_{0}^{i}$. This implies that once $z_{i}(t)$ reaches 00, it remains there forever. Therefore, 00 is an absorbing state. ∎
@@
[…]
@@ B.1 Background on stability of SDEs
${\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
$=\inf_{\Pi\in{\mathcal{C}}}\left(\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|F({\bm{z}})-F(\widetilde{\bm{z}})|\Pi({\mathrm{d}}{\bm{z}},{\mathrm{d}}\widetilde{\bm{z}})\right)$
@@
$\leq L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}))$
$\leq L_{h}c_{1}e^{-c_{2}T}\mathbb{E}\left[(1\wedge|{\bm{x}}-\widetilde{\bm{x}}|)\left(1+V_{2}({\bm{x}})+V_{2}(\widetilde{\bm{x}})\right)\right]$
@@
Now focus on deriving the upper bound for ${\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$. Observe that
${\mathcal{W}}_{2}^{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $={\mathcal{W}}_{2}^{2}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
$=\inf_{\Pi\in{\mathcal{C}}}\left(\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|F({\bm{z}})-F(\widetilde{\bm{z}})|^{2}\Pi({\mathrm{d}}{\bm{z}},{\mathrm{d}}\widetilde{\bm{z}})\right)$
@@
$\leq 2L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}))$
$\leq 2L_{h}c_{1}e^{-c_{2}T}\mathbb{E}\left[(1\wedge|{\bm{x}}-\widetilde{\bm{x}}|)\left(1+V_{2}({\bm{x}})+V_{2}(\widetilde{\bm{x}})\right)\right]$
@@ B.3 Linear Noise SDE and Geometric SDE
${\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
$=\inf_{\Pi\in{\mathcal{C}}}\left(\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|F({\bm{z}})-F(\widetilde{\bm{z}})|\Pi({\mathrm{d}}{\bm{z}},{\mathrm{d}}\widetilde{\bm{z}})\right)$ (24)
@@ D.2 Network Architecture
In practice, we implemented several design choices to optimize the model’s performance and stability. Firstly, we utilized the idea of sinusoidal positional encoding, as proposed by Vaswani et al. (2017), for the time variable, ensuring that each time step possesses a unique encoding. Secondly, we opted for the tanh nonlinearity as the final operation for drift, diffusion, and all other vector fields, following the recommendation by Kidger et al. (2020). This choice is intended to prevent the model from experiencing issues related to excessively large values or gradients. Lastly, we employed layer-wise learning rates for the model’s final layer (e.g. ×$\times 100$), facilitating more precise and adaptive learning for the classification task. This approach can result in improved classification outcomes, as each layer can adjust its learning rate according to the complexity of the features it aims to capture.
@@ D.3 Training strategy
[…]
@@ D.4 Empirical study
Table 8: Classification performance on the ‘BasicMotions’ dataset at a 50% missing rate using different solvers: the explicit Euler, the Milstein, and the Stochastic Runge-Kutta (SRK)(Average taken over five iterations. Runtime, in seconds, for 100 epochs without early-stopping.) Methods Solvers Accuracy Runtime Neural SDE Euler 0.500 ± 0.059 87.4 ± 0.3 Milstein 0.483 ± 0.091 156.9 ± 0.9 SRK 0.500 ± 0.059 749.4 ± 3.4 Neural LSDE Euler 1.000 ± 0.000 89.4 ± 0.5 Milstein 0.983 ± 0.037 112.6 ± 1.0 SRK 1.000 ± 0.000 748.5 ± 2.4 Neural LNSDE Euler 1.000 ± 0.000 95.0 ± 0.4 Milstein 0.983 ± 0.037 135.3 ± 0.8 SRK 0.983 ± 0.037 822.5 ± 2.2 Neural GSDE Euler 0.983 ± 0.037 96.4 ± 0.6 Milstein 0.983 ± 0.037 135.7 ± 1.1 SRK 0.967 ± 0.046 839.1 ± 4.1
Table 9: Classification performance on the ‘BasicMotions’ dataset at a 50% missing rate (Average taken over five iterations. Runtime, in seconds, for 100 epochs without early-stopping.) Methods Accuracy Runtime RNN 0.600 ± 0.070 2.3 ± 0.1 LSTM 0.700 ± 0.095 2.8 ± 0.9 GRU 0.900 ± 0.091 3.0 ± 0.0 GRU-$\Delta t$ 0.967 ± 0.046 34.4 ± 0.3 GRU-D 0.933 ± 0.070 42.4 ± 0.3 MTAN 0.717 ± 0.247 6.2 ± 0.4 MIAM 0.933 ± 0.109 41.7 ± 0.6 GRU-ODE 0.983 ± 0.037 519.8 ± 0.3 ODE-RNN 1.000 ± 0.000 113.3 ± 0.2 ODE-LSTM 0.717 ± 0.126 73.0 ± 0.7 Neural CDE 0.983 ± 0.037 102.5 ± 0.3 Neural RDE 0.717 ± 0.046 93.8 ± 0.4 ANCDE 0.983 ± 0.037 360.5 ± 0.5 EXIT 0.417 ± 0.102 334.5 ± 0.3 LEAP 0.317 ± 0.124 163.0 ± 0.6 Latent SDE 0.350 ± 0.109 238.5 ± 0.6 Neural SDE 0.500 ± 0.059 87.4 ± 0.3 Neural LSDE 1.000 ± 0.000 89.4 ± 0.5 Neural LNSDE 1.000 ± 0.000 95.0 ± 0.4 Neural GSDE 0.983 ± 0.037 96.4 ± 0.6
@@ Appendix E Detailed results of the benchmark datasets
We utilized the Python library torchcde [^12] https://github.com/patrick-kidger/torchcde (Kidger et al., 2020; Morrill et al., 2021b) for the interpolation scheme and the Python library torchsde [^13] https://github.com/google-research/torchsde (Li et al., 2020) to solve Neural SDEs for both the Neural SDE and the proposed methods. This study utilized the original Neural CDE source code [^14] https://github.com/patrick-kidger/NeuralCDE (Kidger et al., 2020) for benchmark methods including GRU-$\Delta t$, GRU-D, GRU-ODE, ODE-RNN, and Neural CDE. Additionally, we employed ODE-LSTM using its original source code [^15] https://github.com/mlech26l/ode-lstms (Lechner & Hasani, 2020), as well as Neural RDE [^16] https://github.com/jambo6/neuralRDEs (Morrill et al., 2021a), ANCDE [^17] https://github.com/sheoyon-jhin/ANCDE (Jhin et al., 2023b), EXIT [^18] https://github.com/sheoyon-jhin/EXIT (Jhin et al., 2022), and LEAP [^19] https://github.com/alflsowl12/LEAP (Jhin et al., 2023a), using their original source code. For Latent SDE, we formulate adjoint SDE proposed by Li et al. (2020) and train backpropagation incorporated with the original KL divergence.

### v2

@@ 2 Related Work & Preliminaries
Let $(\Omega,{\mathcal{F}},P)$ be a probability space. Fix integer $d\geq 1$. For an ${\mathbb{R}}^{d}$-valued random variable $X$, its law is denoted by ${\mathcal{L}}(X)$. $|\cdot|$ is the Euclidean norm where the dimension of the space may vary depending on the context. We denote by $L^{p}(\Omega)$ the set of $X$ with $(\mathbb{E}[|X|^{p}])^{1/p}<\infty$. Let ${\mathcal{P}}({\mathbb{R}}^{d})$ denote the set of probability measures on ${\mathcal{B}}({\mathbb{R}}^{d})$. For two probability measures $\mu,\nu\in{\mathcal{P}}({\mathbb{R}}^{d})$, let ${\mathcal{C}}(\mu,\nu)$ denote the set of probability measures $\Pi$ on ${\mathcal{B}}({\mathbb{R}}^{2d})$ such that its respective marginals are $\mu$ and $\nu$. Then, the Wasserstein distance of order $p\geq 1$ is defined as
@@ 3.1 The proposed Neural SDEs
where the initial condition $h(\cdot;\theta_{h})$ is an affine function with parameter $\theta_{h}$, the drift function $\gamma(\cdot,\cdot;\theta_{\gamma})$ is a neural network with parameter $\theta_{\gamma}$ and the diffusion function $\sigma(\cdot;\theta_{\sigma})$ is a neural network with parameter $\theta_{\sigma}$. Due to its exponential form, Neural GSDE has distinct characteristics associated with deep ReLU networks such as a unique positive solution and an absorbing state of 0, which are not observed in Neural LSDE and Neural LNSDE (see Proposition 3.4 (ii)).
@@ 3.2 Properties of the proposed Neural SDEs
    Assume that the activation function in $\gamma$ is either a tanh or sigmoid function. Then, the SDE in Equation (6) has a unique strong solution ${\bm{z}}(t)$, which is nonnegative almost surely, i.e., $P({\bm{z}}(t)\geq 0\mbox{ for all $t\geq 0$})=1$ a.s. Furthermore, state 0 is an absorbing state.
The proof for Proposition 3.4 can be found in Appendix A. Intuitively, Proposition 3.4 (ii) implies that the latent representation ${\bm{z}}(t)$ of the Neural GSDE has always nonnegative values. Furthermore, once the one of values of the latent representation reaches 0, it remains there forever. This property is consistent with that of ReLU networks which have positive values for activated neurons and zeros for deactivated neurons. In particular, the fact that deactivated neurons never turn to activated neurons corresponds to the property of state 0 being an absorbing state.
@@ 3.3 Robustness under distribution shift
[…]
@@
$)\displaystyle{\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $\leq\sqrt{3}L_{F}L_{h}c_{1}e^{-c_{2}T}\sqrt{\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)}\rho,$
$)\displaystyle{\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $\leq\sqrt{2\sqrt{3}L_{h}c_{1}}L_{F}e^{-c_{2}T/2}\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)^{1/4}\sqrt{\rho},$
@@ Results.
Methods Mean Squared Error ($\times 10^{-3}$)
@@
Table 3: Accuracy and memory usage (in MB) on Speech Commands
@@ Acknowledgement
We thank the teams and individuals for their efforts in dataset preparation and curation for our research, especially the UEA & UCR repository for the 30 datasets we extensively analyzed.
@@
where $\gamma_{i}$ is the $i$-th element of $\gamma$ and $(\sigma(t;\theta_{\sigma}))_{ij}$ represents the $(i,j)$-th element of $\sigma(t;\theta_{\sigma})$, and ${\bm{y}}_{i}(t),{\bm{y}}_{i}(t)]$ is the quadratic variation of ${\bm{y}}_{i}(t)$ up to $t$ for $i=1,2,\ldots,d_{z}$. In the matrix form, ${\bm{z}}(t)$ satisfies
@@
Therefore, ${\bm{z}}(t)$ is the solution of Equation (6). Let $T_{0}^{i}=\inf\{t>0|z_{i}(t)=0\}$ be the first hitting time to state 0 with $z_{i}(0)>0$ for $i=1,2,\ldots,d_{z}$ where $z_{i}(t)$ is the $i$-th component of ${\bm{z}}(t)$. Since ${\mathrm{d}}z_{i}(t)=0$ for $t\geq T_{0}^{i}$, then $z_{i}(t)=0$ for all $t\geq T_{0}^{i}$. This implies that once $z_{i}(t)$ reaches 0, it remains there forever. Therefore, 0 is an absorbing state. ∎
@@
$\mbox{(Neural LSDE)}\quad{\mathrm{d}}{\bm{z}}(t)$ $=\overline{\gamma}_{1}({\bm{z}}(t);\theta_{\gamma},\theta_{\zeta}){\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\mathrm{d}}W(t),$ (12)
$\mbox{(Neural LNSDE)}\quad{\mathrm{d}}{\bm{z}}(t)$ $=\overline{\gamma}_{2}(t,{\bm{z}}(t);\theta_{\gamma},\theta_{\zeta}){\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t){\mathrm{d}}W(t),$ (13)
$\mbox{(Neural GSDE)}\quad\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}$ $=\overline{\gamma}_{2}(t,{\bm{z}}(t);\theta_{\gamma},\theta_{\zeta}){\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\mathrm{d}}W(t).$ (14)
@@ B.1 Background on stability of SDEs
$)\displaystyle{\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
[…]
@@
$)\displaystyle\leq L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}$))
[…]
@@
Now focus on deriving the upper bound for $){\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)). Observe that
$)\displaystyle{\mathcal{W}}_{2}^{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $={\mathcal{W}}_{2}^{2}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
[…]
@@
$)\displaystyle\leq 2L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}$))
[…]
@@ B.3 Linear Noise SDE and Geometric SDE
$)\displaystyle{\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
[…]
@@ D.2 Network Architecture
In practice, we implemented several design choices to optimize the model’s performance and stability. Firstly, we utilized the idea of sinusoidal positional encoding, as proposed by Vaswani et al. (2017), for the time variable, ensuring that each time step possesses a unique encoding. Secondly, we opted for the tanh nonlinearity as the final operation for drift, diffusion, and all other vector fields, following the recommendation by Kidger et al. (2020). This choice is intended to prevent the model from experiencing issues related to excessively large values or gradients. Lastly, we employed layer-wise learning rates for the model’s final layer (e.g. $\times 100$), facilitating more precise and adaptive learning for the classification task. This approach can result in improved classification outcomes, as each layer can adjust its learning rate according to the complexity of the features it aims to capture.
@@ D.3 Training strategy
9:Return $\theta_{\zeta}^{*}$, $\theta_{\gamma}^{*}$, $\theta_{\sigma}^{*}$, $\theta_{\textit{MLP}}^{*}$.
@@ D.4 Empirical study
Table 8: Classification performance on the ‘BasicMotions’ dataset at a 50% missing rate
using different solvers: the explicit Euler, the Milstein, and the Stochastic Runge-Kutta (SRK)
(Average taken over five iterations. Runtime, in seconds, for 100 epochs without early-stopping.)
Methods Solvers Accuracy Runtime
Neural SDE Euler 0.500 ± 0.059 87.4 ± 0.3
Milstein 0.483 ± 0.091 156.9 ± 0.9
SRK 0.500 ± 0.059 749.4 ± 3.4
Neural LSDE Euler 1.000 ± 0.000 89.4 ± 0.5
Milstein 0.983 ± 0.037 112.6 ± 1.0
SRK 1.000 ± 0.000 748.5 ± 2.4
Neural LNSDE Euler 1.000 ± 0.000 95.0 ± 0.4
Milstein 0.983 ± 0.037 135.3 ± 0.8
SRK 0.983 ± 0.037 822.5 ± 2.2
Neural GSDE Euler 0.983 ± 0.037 96.4 ± 0.6
Milstein 0.983 ± 0.037 135.7 ± 1.1
SRK 0.967 ± 0.046 839.1 ± 4.1
Table 9: Classification performance on the ‘BasicMotions’ dataset at a 50% missing rate
(Average taken over five iterations. Runtime, in seconds, for 100 epochs without early-stopping.)
Methods Accuracy Runtime
RNN 0.600 ± 0.070 2.3 ± 0.1
LSTM 0.700 ± 0.095 2.8 ± 0.9
GRU 0.900 ± 0.091 3.0 ± 0.0
GRU-$\Delta t$ 0.967 ± 0.046 34.4 ± 0.3
GRU-D 0.933 ± 0.070 42.4 ± 0.3
MTAN 0.717 ± 0.247 6.2 ± 0.4
MIAM 0.933 ± 0.109 41.7 ± 0.6
GRU-ODE 0.983 ± 0.037 519.8 ± 0.3
ODE-RNN 1.000 ± 0.000 113.3 ± 0.2
ODE-LSTM 0.717 ± 0.126 73.0 ± 0.7
Neural CDE 0.983 ± 0.037 102.5 ± 0.3
Neural RDE 0.717 ± 0.046 93.8 ± 0.4
ANCDE 0.983 ± 0.037 360.5 ± 0.5
EXIT 0.417 ± 0.102 334.5 ± 0.3
LEAP 0.317 ± 0.124 163.0 ± 0.6
Latent SDE 0.350 ± 0.109 238.5 ± 0.6
Neural SDE 0.500 ± 0.059 87.4 ± 0.3
Neural LSDE 1.000 ± 0.000 89.4 ± 0.5
Neural LNSDE 1.000 ± 0.000 95.0 ± 0.4
Neural GSDE 0.983 ± 0.037 96.4 ± 0.6
@@ Appendix E Detailed results of the benchmark datasets
We utilized the Python library torchcde [^12] https://github.com/patrick-kidger/torchcde (Kidger et al., 2020; Morrill et al., 2021b) for the interpolation scheme and the Python library torchsde [^13] https://github.com/google-research/torchsde (Li et al., 2020) to solve Neural SDEs for both the Neural SDE and the proposed methods. This study utilized the original Neural CDE source code [^14] https://github.com/patrick-kidger/NeuralCDE (Kidger et al., 2020) for benchmark methods including $\Delta t, GRU$-D, GRU-ODE, ODE-RNN, and Neural CDE. Additionally, we employed ODE-LSTM using its original source code [^15] https://github.com/mlech26l/ode-lstms (Lechner & Hasani, 2020), as well as Neural RDE [^16] https://github.com/jambo6/neuralRDEs (Morrill et al., 2021a), ANCDE [^17] https://github.com/sheoyon-jhin/ANCDE (Jhin et al., 2023b), EXIT [^18] https://github.com/sheoyon-jhin/EXIT (Jhin et al., 2022), and LEAP [^19] https://github.com/alflsowl12/LEAP (Jhin et al., 2023a), using their original source code. For Latent SDE, we formulate adjoint SDE proposed by Li et al. (2020) and train backpropagation incorporated with the original KL divergence.